OpenStudy (anonymous):

4. If x, y, and z are nonnegative numbers such that x + y = 5 and x + z = 3, find the largest possible value of x^3 + (3/2)y^2 + z^3.

OpenStudy (aum):

\[ \text{Let } S = x^3 + \frac 32y^2 + z^3 \\ x + y = 5; ~~~ y = 5 - x \\ x + z = 3; ~~~ z = 3 - x \\ S = x^3 + \frac 32(5-x)^2 + (3-x)^3 \\ \text{Maximize S.} \\ \]

OpenStudy (anonymous):

Thank you! I was doing some trippy stuff on this problem, haha.

OpenStudy (aum):

But this requires some extra work than a typical maximization problem. dS/dx = 0 yields x = 2 which happens to be a minimum!

OpenStudy (anonymous):

oh noooooooooo

OpenStudy (anonymous):

So we have to check endpt 0 since there's no endpt to the right?

OpenStudy (aum):

"x, y, and z are non-negative numbers" implies: \(x \ge 0\), \(y \ge 0\), \(z \ge 0\) x + y = 5; x = 5 - y Minimum y is 0. So maximum x is 5. x + z = 3; x = 3 - z Minimum z is 0. So maximum x is 3. Therefore, \(0\le x \le 3\). Check S at x = 0 and x = 3 and see which gives max S.

OpenStudy (anonymous):

Oooh okay, thanks, I'll try that right now

OpenStudy (anonymous):

S(0) = 64.5 and S(3) = 33. So max at (0, 64.5)!